# Near-Light Color Photometric Stereo for mono-Chromaticity non-lambertian surface

### v1

@@ Abstract
Color photometric stereo enables single-shot surface reconstruction, extending conventional photometric stereo that requires multiple images of a static scene under varying illumination to dynamic scenarios. However, most existing approaches assume ideal distant lighting and Lambertian reflectance, leaving more practical near-light conditions and non-Lambertian surfaces underexplored. To overcome this limitation, we propose a framework that leverages neural implicit representations for depth and BRDF modeling under the assumption of mono-chromaticity (uniform chromaticity and homogeneous material), which alleviates the inherent ill-posedness of color photometric stereo and allows for detailed surface recovery from just one image. Furthermore, we design a compact optical tactile sensor to validate our approach. Experiments on both synthetic and real-world datasets demonstrate that our method achieves accurate and robust surface reconstruction.
@@ 1 Introduction
Conventional methods, initially proposed by Woodham [1], use least-squares optimization for static Lambertian surfaces and are constrained by distant illumination. Neural network-based approaches [2, 3] have improved performance on non-Lambertian objects. Under near-light illumination setting, LEDPS [4] handles near-light conditions through alternating optimization. FastNFPS [5] has focused on non-Lambertian photometric stereo under uncalibrated near-field lighting. SDMUniPS [6] generalizes the setup to the case of general unknown lighting, enabling reconstruction under more general conditions. However, conventional approaches necessitate capturing multiple images under varying illumination at different timestamps. Since objects are required to be static during the capture, limiting their applications in dynamic scenarios.
@@
Color photometric stereo offers a single-shot alternative through spectral multiplexing, thereby enabling snapshot surface normal estimation. However, for a completely isolated pixel, color photometric stereo is ill-posed. This arises because, in color photometric stereo with $k$ spectral light sources, each isolated pixel provides $k$ observations, but each spectral band has its own distinct spectral reflectance, the total number of unknowns becomes $k+2$ ($k$ spectral reflectances plus two surface normal parameters). To transform this ill-posed problem into a well-posed one, additional priors or constraints are typically required to regularize the solution. Anderson et al. [7, 8] and Y. Ju et al. [9, 10] employed prior information for simplification. Methods [11, 12] simplified the problem by assuming the surface spectral reflectance type to be gray or mono-chromatic with uniform albedo. Guo et al. [13] leveraged multiple pixels sharing the same chromaticity $\mathbf{v}$ to jointly constrain the solution under the assumption of mono-chromaticity. J. Lv et al. [14] addressed the non-Lambertian spectral BRDF under distant light by introducing a strong model assumption of Spectral Reflectance Decomposition, decomposing the BRDF into two simpler components. Physics-Free [15] achieves high-quality reconstruction of non-Lambertian surfaces even under uncalibrated distant illumination. The near-light photometric stereo method of Chen et al. [16] is only applicable to human faces under the Lambertian reflectance assumption. The inherent ill-posedness of color photometric stereo is further exacerbated in practical near-light illumination setting and non-Lambertian reflectances, which has not been well-explored.
@@
To address this problem, we propose a neural implicit modeling framework for near-light color photometric stereo and design a practical capture setup, as shown in Fig. 1. We model the surface depth using the differentiable neural representation, which enables the simultaneous optimization of both depth and normal, thereby reducing the difficulty of surface shape estimation. Simultaneously, under the homogeneous material constraint, the surface reflectance exhibits a unified BRDF structure across the entire surface. This constraint allows the BRDF to be effectively represented using a single implicit neural network. Further incorporating the uniform chromaticity condition, the solution of non-Lambertian spectral BRDF becomes tractable. The integration of these two models transforms the originally ill-posed problem of color photometric stereo with non-Lambertian reflectance under near-field illumination into a well-posed one. This approach facilitates accurate normal estimation and robust geometry reconstruction under near-light condition.
To validate the practical utility of our method, real-world evaluation is necessary. We note that high-resolution tactile sensors such as GelSight [17] operate in a near-field snapshot configuration, which closely matches the operational conditions required by our approach. We identify this as a promising application direction for color photometric stereo. Consequently, we designed a compact near-light visuo-tactile sensor (NearLightTactile) sensor for qualitative evaluation of our method’s performance in real-world scenarios.
@@
  • We propose the first near-light color photometric stereo method, allowing accurate normal recovery for non-Lambertian surface under practical near light setting;
  • We propose to use a neural surface model to jointly recover depth and normal, and a neural BRDF model for mono-chromaticity BRDF, formulating the near-light color photometric stereo to be well-posed;
@@ 2.1 Neural Surface Model
Instead of directly predicting surface normals, we estimate per-pixel depth values and subsequently derive normals analytically, thereby resolving the ill-posedness arising from geometric coupling in near-light color photometric stereo.
@@ 3.1 Experimental Setup
Due to the lack of publicly available color photometric stereo datasets captured under near-light conditions, we generate a synthetic dataset of five objects using Blender v3.6111https://www.blender.org/, designed to cover variations in color, reflectance, and geometry. Our dataset comprises both white and colored objects to examine robustness against intrinsic color; Lambertian and non-Lambertian materials (e.g., metal, ceramic) to test reflectance modeling; and objects ranging from flat to curved and rugged surfaces to evaluate reconstruction across different geometric complexities. All objects are rendered with Blender’s principled BSDF shader. The camera is positioned 35 mm away from the object, with a sensor size of $2.3\times 1.44$ mm, focal length of 2.8 mm, and image resolution of $640\times 480$. A point light source is placed 11 mm away from the camera longitudinally and 21.5 mm laterally. Three colored LEDs (red, green, blue) are simulated to provide near-light illumination, and their intensity is tuned to avoid over-exposure. Rendering is performed in Blender Cycles using a linear color space to preserve radiometric consistency.
Since most color photometric stereo methods require constraints from four or more light sources, we ultimately opted to compare our method with LSPS [1], LEDPS [4], RGB-PS [11] and FastNFPS [5], using their default implementations where available. Performance is evaluated with the standard Mean Angular Error (MAE) in degrees between predicted and ground-truth normals.
@@ 3.2 Experiment Results and Analysis
The quantitative results on the synthetic dataset demonstrate that our method reformulates the originally ill-posed problem of color photometric stereo under near-light illumination into a well-posed one. This is achieved by implicitly disentangling geometry through neural representations while decoupling non-Lambertian spectral BRDF using mono-chromaticity constraints and an implicit BRDF formulation. Through this two-fold decoupling, we show improved reconstruction accuracy under challenging near-light conditions.
@@ 3.4.1 Robustness against Crosstalk
Crosstalk [23] poses a major challenge for color photometric stereo in compact systems , due to spectral overlap in RGB filters and pixel-level cross-coupling in miniaturized sensors. This violates the channel independence assumption in RGB-PS and causes significant artifacts.
@@ 4 Conclusion
In this paper, we presented the first non-Lambertian color photometric stereo method under near-light conditions. By introducing a neural surface model to jointly constrain surface normals and depth, together with a neural BRDF model that enforces a mono-chromaticity constraint, our approach reduces reconstruction ambiguity and makes color photometric stereo tractable. The proposed framework achieves accurate shape recovery from a single multispectral snapshot and demonstrates practical feasibility through real-world validation with a compact visuo-tactile sensor.

### v2

@@ Abstract
Color photometric stereo enables single-shot surface reconstruction, extending conventional photometric stereo that requires multiple images of a static scene under varying illumination to dynamic scenarios. However, most existing approaches assume ideal distant lighting and Lambertian reflectance, leaving more practical near-light conditions and non-Lambertian surfaces underexplored. To overcome this limitation, we propose a framework that leverages neural implicit representations for depth and BRDF modeling under the assumption of mono-chromaticity (uniform chromaticity and homogeneous material), which alleviates the inherent ill-posedness of color photometric stereo and allows for detailed surface recovery from just one image. Furthermore, we design a compact optical tactile sensor to validate our approach. Experiments on both synthetic and real-world datasets demonstrate that our method achieves accurate and robust surface reconstruction. To facilitate future research and ensure reproducibility, we have made our source code publicly available at https://github.com/yulinjiaxue/near-light-color-ps.
@@ 1 Introduction
Conventional methods, initially proposed by Woodham [20], use least-squares optimization for static Lambertian surfaces and are constrained by distant illumination. Neural network-based approaches [9, 4] have improved performance on non-Lambertian objects. Under near-light illumination setting, LEDPS [16] handles near-light conditions through alternating optimization. FastNFPS [12] has focused on non-Lambertian photometric stereo under uncalibrated near-field lighting. SDMUniPS [10] generalizes the setup to the case of general unknown lighting, enabling reconstruction under more general conditions. However, conventional approaches necessitate capturing multiple images under varying illumination at different timestamps.
@@
Color photometric stereo (CPS) offers a single-shot alternative through spectral multiplexing, thereby enabling snapshot surface normal estimation. However, CPS is inherently ill-posed at the pixel level. This arises because, in CPS with $k$ spectral light sources, each isolated pixel provides $k$ observations, but each spectral band has its own distinct spectral reflectance, the total number of unknowns becomes $k+2$ ($k$ spectral reflectances plus two surface normal parameters). To transform this ill-posed problem into a well-posed one, additional priors or constraints are typically required to regularize the solution. Anderson et al. [2] and Y. Ju et al. [11] employed prior information for simplification. Methods [3, 15] simplified the problem by assuming the surface spectral reflectance type to be gray or mono-chromatic with uniform albedo. Guo et al. [6] leveraged multiple pixels sharing the same chromaticity $\mathbf{v}$ to jointly constrain the solution under the assumption of mono-chromaticity. J. Lv et al. [13] addressed the non-Lambertian spectral BRDF under distant light by introducing a strong model assumption of Spectral Reflectance Decomposition, decomposing the BRDF into two simpler components. Physics-Free [8] achieves high-quality reconstruction of non-Lambertian surfaces even under uncalibrated distant illumination. The near-light photometric stereo method of Chen et al. [5] is only applicable to human faces under the Lambertian reflectance assumption. The inherent ill-posedness of CPS is further exacerbated in practical near-light illumination setting and non-Lambertian reflectances.
To address the ill-posedness of CPS under practical near-light illumination and non-Lambertian reflectance, we consider imposing mono-chromaticity prior, which assumes that the surface is uniform chromaticity and homogeneous material. This constraint has been shown to effectively regularize the reconstruction problem under such challenging conditions. Crucially, this assumption is not merely a modeling convenience—it is naturally satisfied in real-world sensing systems such as GelSight [21], which uses a homogeneous elastomer interface under fixed, near-field, single-shot illumination, precisely matching the mono-chromaticity prior. This makes it an ideal application for CPS under this constraint, and motivates the development of tailored photometric stereo methods with both theoretical and practical value.
@@
Under the assumption of mono-chromaticity, we propose a neural implicit modeling framework for near-light CPS and design a practical capture setup, as shown in Fig. 1. We model the surface depth using the differentiable neural representation, which enables the simultaneous optimization of both depth and normal, thereby reducing the difficulty of surface shape estimation. Simultaneously, under the homogeneous material constraint, the surface reflectance exhibits a unified BRDF structure across the entire surface. This constraint allows the BRDF to be effectively represented using a single implicit neural network. Further incorporating the uniform chromaticity condition, the solution of non-Lambertian spectral BRDF becomes tractable. This approach facilitates accurate normal estimation and robust geometry reconstruction under near-light conditions. We also designed a compact near-light visuo-tactile sensor (NearLightTactile) sensor for qualitative evaluation of our method’s performance in real-world scenarios.
@@
  • We propose the first near-light CPS method, allowing accurate normal recovery for non-Lambertian surface under practical near light setting;
  • We propose to use a neural surface model to jointly recover depth and normal, and a neural BRDF model for mono-chromaticity BRDF, formulating the near-light CPS to be well-posed;
@@ 2.1 Neural Surface Model
Instead of directly predicting surface normals, we estimate per-pixel depth values and subsequently derive normals analytically, thereby resolving the ill-posedness arising from geometric coupling in near-light CPS.
@@ 3.1 Experimental Setup
Due to the lack of publicly available CPS datasets captured under near-light conditions, we generate a synthetic dataset of five objects using Blender v3.6111https://www.blender.org/, designed to cover variations in color, reflectance, and geometry. Our dataset comprises both white and colored objects to examine robustness against intrinsic color; Lambertian and non-Lambertian materials (e.g., metal, ceramic) to test reflectance modeling; and objects ranging from flat to curved and rugged surfaces to evaluate reconstruction across different geometric complexities. All objects are rendered with Blender’s principled BSDF shader. The camera is positioned 35 mm away from the object, with a sensor size of $2.3\times 1.44$ mm, focal length of 2.8 mm, and image resolution of $640\times 480$. A point light source is placed 11 mm away from the camera longitudinally and 21.5 mm laterally. Three colored LEDs (red, green, blue) are simulated to provide near-light illumination, and their intensity is tuned to avoid over-exposure. Rendering is performed in Blender Cycles using a linear color space to preserve radiometric consistency.
Since most CPS methods require constraints from four or more light sources, we ultimately opted to compare our method with LSPS [20], LEDPS [16], RGB-PS [3] and FastNFPS [12], using their default implementations where available. Performance is evaluated with the standard Mean Angular Error (MAE) in degrees between predicted and ground-truth normals.
@@ 3.2 Experiment Results and Analysis
The quantitative results on the synthetic dataset demonstrate that our method reformulates the originally ill-posed problem of CPS under near-light illumination into a well-posed one. This is achieved by implicitly disentangling geometry through neural representations while decoupling non-Lambertian spectral BRDF using mono-chromaticity constraints and an implicit BRDF formulation. Through this two-fold decoupling, we show improved reconstruction accuracy under challenging near-light conditions.
@@ 3.4.1 Robustness against Crosstalk
Crosstalk [1] poses a major challenge for CPS in compact systems , due to spectral overlap in RGB filters and pixel-level cross-coupling in miniaturized sensors. This violates the channel independence assumption in RGB-PS and causes significant artifacts.
@@ 4 Conclusion
In this paper, we presented the first non-Lambertian CPS method under near-light conditions. By introducing a neural surface model to jointly constrain surface normals and depth, together with a neural BRDF model that enforces a mono-chromaticity constraint, our approach reduces reconstruction ambiguity and makes CPS tractable. The proposed framework achieves accurate shape recovery from a single multispectral snapshot and demonstrates practical feasibility through real-world validation with a compact visuo-tactile sensor.
5 Acknowledgements
This work was supported by Beijing-Tianjin-Hebei Basic Research Funding Program No. F2024502017, the National Natural Science Foundation of China (Grant No. 62472044, U24B20155, U23B2052, and 52505029), Hebei Natural Science Foundation Project No. 242Q0101Z, Beijing Municipal Science & Technology Program No. Z251100007125021, and Science and Technology Commission of Shanghai Municipality No. 25ZR1401191.